OpenStudy (anonymous):

Find the exact value of the trigonometric function. (If an answer is undefined, enter UNDEFINED.) cos 390°

OpenStudy (freckles):

hint: cos(390)=cos(390-360)

OpenStudy (freckles):

\[\cos(x^o)=\cos(x^o+360^on) \text{ where n is an integer } \\ \sin(x^o)=\sin(x^o+360^on) \text{ where n is an integer }\]

OpenStudy (freckles):

I chose n=-1 since that would put us in the first rotation of a circle